Please help me write the slope-intercept form of the line that passes through the given point and is perpendicular to the graph of each equasion.
(-3,-2),y=x+2

Answer by venugopalramana(3286)

Please help me write the slope-intercept form of the line that passes through the given point and is perpendicular to the graph of each equasion.
(-3,-2),y=x+2
EQN.OF LINE PERPENDICULAR TO THE ABOVE IS OF THE FORM
Y = -X + K
IT IS PASSING THROUGH (-3,-2)..SO..
-2=-(-3)+K=K+3
K=-2-3=-5..HENCE EQN. OF REQUIRED LINE IS
Y = -X - 5